OpenStudy (anonymous):

LIM X TEND S TO INFINITY 3X*X+5X+4\X*X*X+7X

OpenStudy (anonymous):

0

OpenStudy (anonymous):

Note: the degree in the numerator is 2 and the degree in the denominator is 3 x^2/x^3 simplifies to 1/x as 1/x gets bigger, we get 0

OpenStudy (anonymous):

\[\Large \mathop {\lim }\limits_{x \to \infty } \frac{{3{x^2} + 5x + 4}}{{{x^3} + 7x}} = \mathop {\lim }\limits_{x \to \infty } \frac{{3{x^2}}}{{{x^3}}} = 0\]

OpenStudy (anonymous):

Because the order of the denominator is greater than the order of the numerator, the limit will be zero because the denominator grows faster than the numerator. 1/infinity equals zero